OpenStudy (anonymous):

Kathryn currently has an account balance of $2,248.95. She opened the account 7 years ago with a deposit of $1,762.14. If the interest compounds quarterly, what is the interest rate on the account?

OpenStudy (anonymous):

" interest compounds quarterly" Same formula with i = r/4 and nt = 4 t

OpenStudy (anonymous):

Yeah but in this one I have to look for the Interest rate

OpenStudy (anonymous):

$2,248.95/ $1,762.14 = ( 1 + r/4 ) ^28 Exponent both by 1/28:

OpenStudy (anonymous):

* both sides

OpenStudy (anonymous):

3.5%

OpenStudy (anonymous):

Does the rate make sense to you?